OpenStudy (anonymous):

Find the sum of the first 9 terms of the sequence. Show all work for full credit. 2, -7, -16, -25, ...

OpenStudy (anonymous):

Are you familiar with Sigma notation?

OpenStudy (anonymous):

not at all

OpenStudy (texaschic101):

you could just use the two formulas...one to find the 9th term and one to find the sum

OpenStudy (texaschic101):

you can read about it here.. http://www.regentsprep.org/Regents/math/algtrig/ATP2/ArithSeq.htm or just wait and see what happytales says

OpenStudy (anonymous):

OMG

OpenStudy (anonymous):

IT JUST BACKSPACED.... OMGGG ARE YOU KIDDING ME

OpenStudy (anonymous):

Stupid browser shortcuts -_- let me restart....

OpenStudy (anonymous):

Sigma notation is dead useless for solving arithmetic and geometric sequences (and also applications in calculus). It is typically the most popular method in solving sequences. The symbol sigma means "to sum". It looks like this: \[\sum_{n=1}^{x}\] The value at the bottom of the symbol indicates the starting value while the top value indicates how many terms to sum up. For this arithmetic sequence, we can use equation: \[\sum_{k=0}^{n-1}(a+kd)=\frac{ n }{ 2}(2a+(n-1)d)\] a= First term d=Common difference n=Number of terms

OpenStudy (anonymous):

Useful*